# A model-free approach to control barrier functions using funnel control

### v1

@@ Abstract
Keywords: control barrier functions, funnel control, model-free control, uncertain systems, nonlinear output-feedback
@@ I Introduction
Control barrier functions (CBFs) can ensure satisfaction of safety-critical constraints and they have gained popularity in robotics applications because of their simplicity and modularity, see [4, 5] for an overview. Specifically, CBFs determine the control input by solving a quadratic program (QP) that enforces a desired bound on the derivative of the barrier, which then ensures positive invariance and asymptotic stability of a safe set. However, application of CBF-based control requires a model that describes the nonlinear dynamics, measurement of the state, and implementation complexity scales with the dimension of the nonlinear model. The first issues have been addressed in the literature by investigating robustness to bounded disturbances [6, 7], sector nonlinearities [8], or state estimation error [9], and by analyzing input-to-state safety [10, 11]. Furthermore, some approaches investigate extracting the required model information from data [12, 13], which, however, also increases the complexity of application. To mitigate complexity, [14] investigate the utilization of simplified reduced-order models. Furthermore, for a special class of reduced-order kinematic models, [15] provides a model-free CBF-based control law, which requires minimal knowledge about the system dynamics. In this paper, we develop a new model-free CBF-based approach by leveraging funnel control.
@@ II System class and control objective
    the internal dynamics are bounded-input bounded-state stable, i.e., for all $c_{0}>0$ there exists $\bar{q}>0$ such that for all $\eta^{0}\in\mathbb{R}^{n-m}$ and all $\zeta\in L^{\infty}(\mathbb{R}_{\geq 0},\mathbb{R}^{m})$ it holds that $\|\eta^{0}\|+\|\zeta\|_{\infty}\leq c_{0}$ implies ${\|\eta(\cdot;0,\eta^{0},\zeta)\|_{\infty}\leq\bar{q}}$, where $\eta(\cdot;0,\eta^{0},\zeta):\mathbb{R}_{\geq 0}\to\mathbb{R}^{n-m}$ denotes the unique global solution of the second of equations (2) for ${y=\zeta}$ [^1] The aforementioned assumption ensures that the maximal solution $\eta(\cdot;0,\eta^{0},\zeta)$ can be extended to a global solution, cf. [23, § 10, Thm. XX].
Assumption 1a) means in particular that the systems under consideration have global relative degree one, cf. [24, Def. 5.1]. Although we do not assume symmetry, we note that for many mechanical systems, $(L_{G}H)(x)$ simplifies to an inertia-related matrix, which is inherently symmetric and positive-definite and thus satisfies Assumption 1a), cf. [25]. While assuming a well-defined relative degree is common in high-gain adaptive control, cf. [16, 26], we restrict ourselves to systems with a relative degree one for simplicity of presentation of the main ideas. A generalization of the results presented here will be topic of future work. Instead of positive definiteness of $g(\cdot)$, negative definiteness is a viable assumption as well, which will only change the sign in the feedback law derived later. We emphasize that, while existence of the diffeomorphism $T$ putting system (1) in the form (2) is assumed in Assumption 1b), concrete knowledge about this transformation is not required – it is merely used as a tool in the following analysis. Sufficient conditions for its existence can be found in [24, Cor. 5.6]. Assumption 1c) ensures that the unmeasured internal state evolves within a (unknown) compact set for bounded $y$. In what follows, we call a (locally) absolutely continuous function $x:[0,\omega)\to\mathbb{R}^{n}$, $\omega\in(0,\infty]$, with $x(0)=x^{0}$ a solution (in the sense of Carathéodory) to (1), if it satisfies (1) for almost all $t\in[0,\omega)$. A solution $x$ is said to be maximal, if it has no right extension that is also a solution; a maximal solution is global, if $\omega=\infty$.
@@ III-A Control barrier functions
Following [5, Def. 2], we recall the following concepts.
@@
The following proposition characterizes the safety properties of CBF-based controllers and is based on [5, Thm. 2]. [^3] The strict inequality in Def. 1 ensures $\big{(}\partial_{t}b(t,x),\nabla_{x}b(t,x)\big{)}\neq 0$ for all $(t,x)\in\mathcal{D}$ with $b(t,x)=0$ and that the set $K_{\mathrm{CBF}}(t,x,\alpha)$ is non-empty for all $(t,x)\in\mathcal{D}$, see also [14, Rem. 1].
@@ III-B Main result
We note that $b$ is continuously differentiable on $\mathcal{D}$. For the sake of better legibility, we omit the arguments of all functions whenever it is clear from the context. We use the notation $e=H(x)-y_{r}$ and the diffeomorphism $T$ as well as the functions in (2) from Assumption 1. Inserting $u=(g\circ T)^{-1}(x)\cdot(\max\{2c,1\}e-(f\circ T)(x)+\dot{y}_{r})$, which we are allowed to select with $c>0$ from (5), we obtain
@@
$=\dot{\psi}\psi+\langle e,-\dot{y}_{r}+f\circ T\rangle$
$\quad+\langle e,(g\circ T)(g\circ T)^{-1}(\max\{2c,1\}e-f\circ T+\dot{y}_{r})\rangle$
@@
with user-chosen constants $\overline{k}\geq\underline{k}>0$, where we recall that $b(t,y)>0$ for $(t,y)\in\text{int}(\mathcal{C})$. Note that no system data is used to define the set $\mathcal{U}(t,y)$. The feedback given by (8) resembles a funnel controller, cf. [16, Eq. (2)]. The following result shows that control laws based on the model-free set $\mathcal{U}(t,y)$ can be interpreted as CBF-based controls.
@@
The existence of a local maximal solution ${x:[0,\omega)\to\mathbb{R}^{n}}$, $\omega\in(0,\infty]$, follows from [16, Thm. 5]. We show statement (ii) with $\varepsilon\in(0,1)$ determined below. Note that for any $\zeta\in C(\mathbb{R}_{\geq 0},\mathbb{R}^{m})$ with $\|\zeta\|_{\infty}\leq\bar{\psi}+\bar{y}_{r}$ it follows from Assumption 1 that the global solution $\eta(\cdot;\eta^{0},\zeta)$ exists. Then, again by Assumption 1, there exists $\bar{q}>0$ such that $\|\eta(\cdot;\eta^{0},\zeta)\|_{\infty}\leq\bar{q}$ for any such function $\zeta$. We set
@@
However, to ensure asymptotic stability of $\mathcal{C}$ (cf. [4, Prop. 2]), the constant $\underline{k}/\delta>0$ needs to be chosen sufficiently large, i.e., system knowledge has to be included in the controller design. Nevertheless, this sets the proposed controller design apart from funnel control approaches [16, 17, 18] as they cannot achieve the evolution of the tracking error $e$ within the funnel boundaries at a later point in time, if the initial error $e(0)$ does not satisfy $(0,y(0))\in\mathrm{int}(\mathcal{C})$.
@@ IV-A Optimization-based CBF implementation
Here, $\mathrm{cost}(t,u,y)$ is a convex quadratic cost function while the set $\mathcal{U}(t,y)$ from (8) can be written as a set of linear inequality constraints, i.e., (10) is a QP. The quadratic cost can encode a wide range of different performance criteria, see [5]. Of particular relevance is penalizing the difference to a desired input, i.e., $\mathrm{cost}(t,u,y)=\|u-u_{r}(t)\|^{2}$, which is also referred to as a minimally invasive safety filter [1]. In this case, Problem (10) has an analytical solution (cf. [14, Eq. (15)]). We note that there exist results ensuring Lipschitz continuity of the resulting feedback $\mu(t,y)$, cf. [6]. However, due to a different structure of $\mathcal{U}(t,y)$, it is not obvious if these results also apply for the proposed implementation (10). This requires further investigation.
@@ IV-D Prescribed performance control and reciprocal CBF
As mentioned in Section I, a connection between PPC and controls based on reciprocal CBFs has been discovered in [22]. By setting the auxiliary variables $e_{1}(t,x):=(x-x_{r}(t))/\psi(t)$ and $\xi(t,x):=e_{1}(t,x)/\sqrt{1-\|e_{1}(t,x)\|^{2}}$, one may define ${B_{\rm FC}:\mathbb{R}^{n}\to\mathbb{R}_{\geq 0}}$, $\xi\mapsto\tfrac{1}{2}\|\xi\|^{2}$. Using $B_{\rm FC}$, we can define the feedback $u(t,x)=-k\nabla_{x}B_{\rm FC}(\xi(t,x))$ for some $k>0$. Then, thanks to the results in [22], we can show that i) $u$ keeps the system within predefined boundaries, i.e., $\|x(t)-x_{r}(t)\|<\psi(t)$ (cf. [22, Thm. 2]), ii) $B_{\rm FC}$ is a reciprocal CBF (RCBF) (cf. [22, Thm. 2]), and iii) $u$ defines a RCBF-based controller (cf. [22, Thm. 4]). In comparison, key benefits of the approach proposed in the present paper are the fact that zeroing CBFs are used, internal dynamics are considered and, most importantly, the implementation can be naturally adjusted to take different optimization criteria into account (cf. Section IV-A) and provide a control strategy to return to the safe set, cf. (9).

### v2

@@ Abstract
Keywords: Control Barrier Functions, Adaptive control, Nonlinear output feedback, Uncertain systems,
@@ I Introduction
Control barrier functions (CBFs), first introduced in [4], can ensure satisfaction of safety-critical constraints and they have gained popularity in robotics applications because of their simplicity and modularity, see [5, 6] for an overview. Specifically, CBFs determine the control input by solving a quadratic program (QP) that enforces a desired bound on the derivative of the barrier, which then ensures positive invariance and asymptotic stability of a safe set. However, application of CBF-based control requires a model that describes the nonlinear dynamics, measurement of the state, and implementation complexity scales with the dimension of the nonlinear model. The first issues have been addressed in the literature by investigating robustness to bounded disturbances [7, 8], sector nonlinearities [9], or state estimation error [10], and by analyzing input-to-state safety [11, 12]. Furthermore, some approaches investigate extracting the required model information from data [13, 14], which, however, also increases the complexity of application. To mitigate complexity, [15] investigate the utilization of simplified reduced-order models. Furthermore, for a special class of reduced-order kinematic models, [16] provides a model-free CBF-based control law, which requires minimal knowledge about the system dynamics. In this paper, we develop a new model-free CBF-based approach by leveraging funnel control.
@@ II System class and control objective
    the internal dynamics are bounded-input bounded-state stable, i.e., for all $c_{0}>0$ there exists $\bar{q}>0$ such that for all $\eta^{0}\in\mathbb{R}^{n-m}$ and all $\zeta\in L^{\infty}(\mathbb{R}_{\geq 0},\mathbb{R}^{m})$ it holds that $\|\eta^{0}\|+\|\zeta\|_{\infty}\leq c_{0}$ implies ${\|\eta(\cdot;0,\eta^{0},\zeta)\|_{\infty}\leq\bar{q}}$, where $\eta(\cdot;0,\eta^{0},\zeta):\mathbb{R}_{\geq 0}\to\mathbb{R}^{n-m}$ denotes the unique global solution of the second of equations (2) for ${y=\zeta}$ [^1] The aforementioned assumption ensures that the maximal solution $\eta(\cdot;0,\eta^{0},\zeta)$ can be extended to a global solution, cf. [24, § 10, Thm. XX].
Assumption 1a) means in particular that the systems under consideration have global relative degree one, cf. [25, Def. 5.1]. Although we do not assume symmetry, we note that for many mechanical systems, $(L_{G}H)(x)$ simplifies to an inertia-related matrix, which is inherently symmetric and positive-definite and thus satisfies Assumption 1a), cf. [26]. While assuming a well-defined relative degree is common in high-gain adaptive control, cf. [17, 27], we restrict ourselves to systems with a relative degree one for simplicity of presentation of the main ideas. A generalization of the results presented here will be topic of future work. Instead of positive definiteness of $g(\cdot)$, negative definiteness is a viable assumption as well, which will only change the sign in the feedback law derived later. We emphasize that, while existence of the diffeomorphism $T$ putting system (1) in the form (2) is assumed in Assumption 1b), concrete knowledge about this transformation is not required – it is merely used as a tool in the following analysis. Sufficient conditions for its existence can be found in [25, Cor. 5.6]. Assumption 1c) ensures that the unmeasured internal state evolves within a (unknown) compact set for bounded $y$. In what follows, we call a (locally) absolutely continuous function $x:[0,\omega)\to\mathbb{R}^{n}$, $\omega\in(0,\infty]$, with $x(0)=x^{0}$ a solution (in the sense of Carathéodory) to (1), if it satisfies (1) for almost all $t\in[0,\omega)$. A solution $x$ is said to be maximal, if it has no right extension that is also a solution; a maximal solution is global, if $\omega=\infty$.
@@ III-A Control barrier functions
Following [6, Def. 2], we recall the following concepts.
@@
The following proposition characterizes the safety properties of CBF-based controllers and is based on [6, Thm. 2]. [^3] The strict inequality in Def. 1 ensures $\big{(}\partial_{t}b(t,x),\nabla_{x}b(t,x)\big{)}\neq 0$ for all $(t,x)\in\mathcal{D}$ with $b(t,x)=0$ and that the set $K_{\mathrm{CBF}}(t,x,\alpha)$ is non-empty for all $(t,x)\in\mathcal{D}$, see also [15, Rem. 1].
@@ III-B Main result
We note that $b$ is continuously differentiable on $\mathcal{D}$. For the sake of better legibility, we omit the arguments of all functions whenever it is clear from the context. We use the notation $e=H(x)-y_{r}$ and the diffeomorphism $T$ as well as the functions in (2) from Assumption 1. Inserting the input $u=(g\circ T)^{-1}(x)\cdot(-\max\{2c,1\}e-(f\circ T)(x)+\dot{y}_{r})$, which we are allowed to select with $c>0$ from (5), we obtain
@@
$=\dot{\psi}\psi-\langle e,-\dot{y}_{r}+f\circ T\rangle$
$\quad-\langle e,(g\circ T)(g\circ T)^{-1}(-\max\{2c,1\}e-f\circ T+\dot{y}_{r})\rangle$
@@
with user-chosen constants $\overline{k}\geq\underline{k}>0$, where we recall that $b(t,y)>0$ for $(t,y)\in\text{int}(\mathcal{C})$. Note that no system data is used to define the set $\mathcal{U}(t,y)$. The feedback given by (8) resembles a funnel controller, cf. [17, Eq. (2)]. The following result shows that control laws based on the model-free set $\mathcal{U}(t,y)$ can be interpreted as CBF-based controls.
@@
The existence of a local maximal solution ${x:[0,\omega)\to\mathbb{R}^{n}}$, $\omega\in(0,\infty]$, follows from [17, Thm. 5]. We show statement (ii) with $\varepsilon\in(0,1)$ determined below. Note that for any $\zeta\in C(\mathbb{R}_{\geq 0},\mathbb{R}^{m})$ with $\|\zeta\|_{\infty}\leq\bar{\psi}+\bar{y}_{r}$ it follows from Assumption 1 that the global solution $\eta(\cdot;\eta^{0},\zeta)$ exists. Then, again by Assumption 1, there exists $\bar{q}>0$ such that $\|\eta(\cdot;\eta^{0},\zeta)\|_{\infty}\leq\bar{q}$ for any such function $\zeta$. We set
@@
However, to ensure asymptotic stability of $\mathcal{C}$ (cf. [5, Prop. 2]), the constant $\underline{k}/\delta>0$ needs to be chosen sufficiently large, i.e., system knowledge has to be included in the controller design. Nevertheless, this sets the proposed controller design apart from funnel control approaches [17, 18, 19] as they cannot achieve the evolution of the tracking error $e$ within the funnel boundaries at a later point in time, if the initial error $e(0)$ does not satisfy $(0,y(0))\in\mathrm{int}(\mathcal{C})$.
@@ IV-A Optimization-based CBF implementation
Here, $\mathrm{cost}(t,u,y)$ is a convex quadratic cost function while the set $\mathcal{U}(t,y)$ from (8) can be written as a set of linear inequality constraints, i.e., (10) is a QP. The quadratic cost can encode a wide range of different performance criteria, see [6]. Of particular relevance is penalizing the difference to a desired input, i.e., $\mathrm{cost}(t,u,y)=\|u-u_{r}(t)\|^{2}$, which is also referred to as a minimally invasive safety filter [1]. In this case, Problem (10) has an analytical solution (cf. [15, Eq. (15)]). We note that there exist results ensuring Lipschitz continuity of the resulting feedback $\mu(t,y)$, cf. [7]. However, due to a different structure of $\mathcal{U}(t,y)$, it is not obvious if these results also apply for the proposed implementation (10). This requires further investigation.
@@ IV-D Prescribed performance control and reciprocal CBF
As mentioned in Section I, a connection between PPC and controls based on reciprocal CBFs has been discovered in [23]. By setting the auxiliary variables $e_{1}(t,x):=(x-x_{r}(t))/\psi(t)$ and $\xi(t,x):=e_{1}(t,x)/\sqrt{1-\|e_{1}(t,x)\|^{2}}$, one may define ${B_{\rm FC}:\mathbb{R}^{n}\to\mathbb{R}_{\geq 0}}$, $\xi\mapsto\tfrac{1}{2}\|\xi\|^{2}$. Using $B_{\rm FC}$, we can define the feedback $u(t,x)=-k\nabla_{x}B_{\rm FC}(\xi(t,x))$ for some $k>0$. Then, thanks to the results in [23], we can show that i) $u$ keeps the system within predefined boundaries, i.e., $\|x(t)-x_{r}(t)\|<\psi(t)$ (cf. [23, Thm. 2]), ii) $B_{\rm FC}$ is a reciprocal CBF (RCBF) (cf. [23, Thm. 2]), and iii) $u$ defines a RCBF-based controller (cf. [23, Thm. 4]). In comparison, key benefits of the approach proposed in the present paper are the fact that zeroing CBFs are used, internal dynamics are considered and, most importantly, the implementation can be naturally adjusted to take different optimization criteria into account (cf. Section IV-A) and provide a control strategy to return to the safe set, cf. (9).